OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

if a number is solution for an equation it means that the number satisfies the given condition consider an eqn x - y = 1 x + y = 2 this both eqns contains variables x and y solution for this eqn is x= 3/2 and y = 1/2

OpenStudy (anonymous):

4x=0 is an example, when we talk about solution it means that we want to find all numbers that when you make them equal x, the equation is still true. Then for this one(a very simple one) the solution would be x=0. For more complicated equations like x^2+2=6 you need to isolate x by multiplying and summing on both sides like this: x^2+2=6-->x^2+2-2=6-2-->x^2=4 (now take the square root on both sides)-->x=+-2, and the +- is because the negative number and the positive number both give the same thing when squared.